What is the answer to x^2+6*x=16
I got to the point where I moved the "16" to the left and I make the whole equation zero and then I have no clue what to do next. I know that the answer is (2,-8) but I want to know how to get the answer.

Thanks for your help.

Answer by richwmiller(17219)

If you know the answer then solving it is easy
x=2 x=-8
(x-2)*(x+8)=0
FOIL
x^2+8x-2x-16=0
x^2+6x-16=0
x^2+6x=16
now going in reverse
x^2+6x=16
x^2+6x-16=0
we want factors of -16 which add up to 6
-2 and 8 come to mind
regroup them
x^2-2x+8x-16=0
x*(x-2)+8*(x-2)=0
(x+8)*(x-2)=0
x=-8 x=2